How much money would you have to deposit today in order to have $3,000 in four years if the discount rate is 8 percent per year? (Do not round intermediate calculations. Round your answer to 2 decimal places. (e.g., 32.16))
Answer #1

The amount of money that needs to be deposited today is :

PV ( 1 + r)^n = FV

PV * (1.08)^4 = $3000

PV = $3,000/ 1.3605

= $2205.0717

= $2205.07 ( ROUNDED OFF TO TWO DECIMAL PLACES)

Alternatively,

FV = $3000

N = 4 YEARS

I/Y = 8%

COMPUTE PV

PV = $2205.09 ( ROUNDED OFF TWO DECIMAL PLACES)